OpenStudy (setsuna-yuregeshi):

I need help to answer an equation. I don't just want the answer, but how to do it...

OpenStudy (setsuna-yuregeshi):

The equation is: \[\frac{ 1 }{x } + \frac{ 1 }{x+1} = \frac{ 17 }{72}\]

OpenStudy (setsuna-yuregeshi):

@phi

OpenStudy (setsuna-yuregeshi):

Anyone know what I need to do???? PlEASE HELP

OpenStudy (setsuna-yuregeshi):

@DimamndHeart15713

OpenStudy (setsuna-yuregeshi):

@TheSmartOne

OpenStudy (setsuna-yuregeshi):

@Luigi0210

OpenStudy (setsuna-yuregeshi):

@SarthKohli

OpenStudy (setsuna-yuregeshi):

@Astrophysics

TheSmartOne (thesmartone):

You need to add the two fractions. Find the LCM of x and x + 1

OpenStudy (setsuna-yuregeshi):

ok, yeah So the LCM is x+1 right?

OpenStudy (phi):

one way is to "clear the denominator" multiply both sides (and all terms) by x(x+1)

OpenStudy (setsuna-yuregeshi):

@phi I haven't learned that yet. I need to do the LCM thing..

OpenStudy (phi):

it's the same idea. if you do it in steps: multiply both sides (and all terms) by x

OpenStudy (setsuna-yuregeshi):

So, what's the LCM of this?

OpenStudy (setsuna-yuregeshi):

x+1?

OpenStudy (phi):

you would get \[ \frac{ x }{x } + \frac{ x }{x+1} = \frac{ 17x }{72}\\ 1+ \frac{ x }{x+1} = \frac{ 17x }{72} \]

OpenStudy (setsuna-yuregeshi):

OOh ok,

OpenStudy (phi):

now multiply by (x+1) both sides, all terms.

OpenStudy (setsuna-yuregeshi):

So,

OpenStudy (phi):

you would get \[ x+1 + x\frac{x+1}{x+1} = \frac{17x(x+1)}{72} \\ x+1 + x= \frac{17x(x+1)}{72} \]

OpenStudy (setsuna-yuregeshi):